# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v2

@@ 2 Notations, Assumptions and Measures
Throughout this paper, we characterize the influence of the mixing matrices using only $s_{A}$, $s_{B}$, $c$, and $n$. A connection between these quantities and the previously discussed metrics is established in Proposition 2.
Proposition 2 (Relation with other measures).
@@
where $M_{A}:=\max_{i\geq 0}\|A^{i}-A_{\infty}\|_{2}$, $M_{B}:=\max_{i\geq 0}\|B^{i}-B_{\infty}\|_{2}$.
@@
3 Preliminary
In this section, we present a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms. We also discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
3.1 Push-Only and Pull-Only methods align with centralized SGD
Most decentralized algorithms—such as decentralized SGD [36], gradient tracking [31, 32, 43, 44], gradient-Push [45], Push-DiGing [7], gradient-Pull [9], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting. These algorithms can generally be transformed into the following form:
$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-\alpha\bar{g}^{(k)}+\,\epsilon^{(k)}$ (5)
where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}(x_{i}^{(k)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local optimization variables, with $v_{i}^{(k)}$ denoting the combination weights, and $\bar{g}^{(k)}=\frac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$ is the average stochastic gradient. The term $\epsilon^{(k)}$ captures additional error due to the decentralized nature of the algorithm.
For Push-Only and Pull-Only optimization algorithms, the additional error term $\epsilon$ is carefully designed to be either zero or to decay exponentially fast. This design ensures that the trajectories of the local decision variables closely track those of centralized SGD. When the error is sufficiently small, the linear speedup result naturally follows from the same analytical framework used for centralized SGD. Below, we provide a brief example illustrating how to derive the formulation in (5).
@@ 3.1 Push-Only and Pull-Only methods align with centralized SGD
The Push-Diging algorithm iterates as follows:
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A_{k}{\mathbf{x}}^{(k)}-\alpha V_{k+1}^{-1}{\mathbf{y}}^{(k)},$ (6a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (6b)
where $v_{k}=B^{k}\mathds{1}$, $V_{k}=\mathrm{Diag}(v_{k})$ is a diagonal matrix, $A_{k}=V_{k+1}^{-1}BV_{k}$, and ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient defined in the notation section. By left-multiplying both sides of (6a) by $n^{-1}v_{k+1}^{\top}$, we obtain
$n^{-1}v_{k+1}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}.$
Letting $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}$, this update exactly matches the form in (5).
@@
FROST/PullDiag-GT algorithm only relies on a row-stochastic matrix $A$. Its recursion can be written as follows:
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (7a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=A{\mathbf{y}}^{(k)}+D_{k+1}^{-1}{\mathbf{g}}^{(k+1)}-D_{k}^{-1}{\mathbf{g}}^{(k)},$ (7b)
@@
where $D_{k}=\mathrm{Diag}(A^{k})$ is a diagonal matrix, ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. By left-multiplying both sides of (7a) by $n^{-1}\pi_{A}^{\top}$, we obtain
$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}+\epsilon^{(k)},$
@@
where $\epsilon^{(k)}=\alpha n^{-1}\left(\mathds{1}^{\top}-\pi_{A}^{\top}D_{k}^{-1}\right){\mathbf{g}}^{(k)}$ converges to zero exponentially fast [Lemma 12, [39]]. Letting $\hat{x}^{(k)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, this update matches the form  (5).
Figure 1: The extra error does not vanish. The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant. The experiment is a logistic regression on synthetic dataset. We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005, and detailed settings is referred to Appendix C.
3.2 Push-Pull suffers from extra error terms
Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (8a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (8b)
where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the topology, $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$, ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying both sides of (8b) by $\mathds{1}^{\top}$, we achieve
$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(k+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(k)}+\mathds{1}^{\top}{\mathbf{g}}^{(k+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k+1)},$ (9)
where the last equality holds due to the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Next, we left-multiply both sides of (8a) by $\pi_{A}^{\top}$, which yields
@@ 3.2 Push-Pull suffers from extra error terms
$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}$
$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}.$
Let $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, and recall from (9) that $\mathds{1}^{\top}{\mathbf{y}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k)}$. Then, we have
@@
$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$ (10)
where $c=n\pi_{A}^{\top}\pi_{B}$ is a constant and $\bar{g}^{(k)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$. We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\|\bar{g}^{(k)}\|^{2}$ is the dominating factor. But here, both the squared error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|^{2}$ and the cross product $\left\langle\bar{g}^{(k)},\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\right\rangle$ can introduce additional noise that may dominate the noise in $\bar{g}^{(k)}$.
For certain specific topologies, the error term can be ignored or absorbed. For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull (BTPP) and demonstrates the linear speedup result on this particular topology. In BTPP, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, leading to $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$. By setting the loss function at the first node to zero, $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$ is achieved, which simplifies the process of achieving linear speedup. Another case is when $A$ is doubly-stochastic; we have $\pi_{A}^{\top}(I-B_{\infty})=0$, which removes the error term.
Unfortunately, for general topologies, the error term lacks exploitable structure. This usually leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result. Therefore, we recognize the central challenge in the analysis of stochastic Push-Pull: How to estimate the noise level in error terms more accurately? We tackle this key question in the following section.

### v3

@@ 2 Notations, Algorithms, Assumptions and Metrics
By the exponential decay property (3), both $s_{A}$ and $s_{B}$ are finite. In this paper, the influence of the mixing matrices is described solely through $s_{A}$, $s_{B}$, $c$, and $n$. Their connection to previously introduced metrics is given in the next proposition.
Proposition 2 (Relation with existing metrics).
@@
where $M_{A}:=\max_{t\geq 0}\|A^{t}-A_{\infty}\|_{2}$ and $M_{B}:=\max_{t\geq 0}\|B^{t}-B_{\infty}\|_{2}$ are finite numbers which can be proved to be no larger than $\sqrt{n}$.
@@
3 Intrinsic Differences between Push-Only, Pull-Only, and Push-Pull
This section provides a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms, clarifying their key differences from Push-Pull. We also highlight the limitations of existing analyses and explain why they cannot establish the linear speedup property of Push-Pull.
3.1 Push-Only and Pull-Only align with centralized SGD
Many decentralized algorithms—such as decentralized SGD [36], gradient tracking [32, 33, 34], Push-Diging [9], Gradient-Pull [11], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting. These algorithms can be reformulated as
$\displaystyle\hat{x}^{(t+1)}=\underbrace{\hat{x}^{(t)}-\alpha\bar{g}^{(t)}}_{\text{centralized SGD}}+\epsilon^{(t)},$ (7)
where $\hat{x}^{(t)}=\sum_{i=1}^{n}v_{i}^{(t)}(x_{i}^{(t)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local decision variables $\{x_{i}^{(t)}\}_{i=1}^{n}$, with $v_{i}^{(t)}$ denoting the combination weights. The term $\bar{g}^{(t)}=\tfrac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(t)}$ is the globally averaged stochastic gradient, and $\epsilon^{(t)}$ captures the residual error due to decentralization. When $\epsilon^{(t)}=0$, the recursion in (7) reduces to centralized SGD, whose linear speedup property is well established (see Table 1).
For Push-Only and Pull-Only algorithms, the residual $\epsilon^{(t)}$ is either exactly zero or decays exponentially to zero. This ensures that the trajectory of the weighted average of local iterates closely tracks that of centralized SGD. Once the residual is sufficiently small, the linear speedup property follows directly from the standard analysis of centralized SGD. We illustrate this alignment with two examples.
@@ 3.1 Push-Only and Pull-Only align with centralized SGD
Push-DiGing is a Push-Only method, whose iterates are given by
$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A_{t}{\mathbf{x}}^{(t)}-\alpha V_{t+1}^{-1}{\mathbf{y}}^{(t)},$ (8a)
$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=B{\mathbf{y}}^{(t)}+{\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)},$ (8b)
where $B$ is column-stochastic, $v_{t}=B^{t}\mathds{1}$, $V_{t}=\mathrm{Diag}(v_{t})$, $A_{t}=V_{t+1}^{-1}BV_{t}$, and ${\mathbf{g}}^{(t)}$ stacks local stochastic gradients. Left-multiplying (8a) by $n^{-1}v_{t+1}^{\top}$ yields
$n^{-1}v_{t+1}^{\top}{\mathbf{x}}^{(t+1)}=n^{-1}v_{t}^{\top}{\mathbf{x}}^{(t)}-\alpha\bar{g}^{(t)}.$
Defining $\hat{x}^{(t)}=n^{-1}v_{t}^{\top}{\mathbf{x}}^{(t)}$, the recursion matches (7) with $\epsilon^{(t)}=0$.
@@
FROST/PullDiag-GT is a Pull-Only method based on a row-stochastic matrix $A$. Its updates are
$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A{\mathbf{x}}^{(t)}-\alpha{\mathbf{y}}^{(t)},$ (9a)
$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=A{\mathbf{y}}^{(t)}+D_{t+1}^{-1}{\mathbf{g}}^{(t+1)}-D_{t}^{-1}{\mathbf{g}}^{(t)},$ (9b)
@@
where $D_{t}=\mathrm{Diag}(A^{t})$, ${\mathbf{g}}^{(t)}$ is the stacked stochastic gradient, and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying (9a) by $n^{-1}\pi_{A}^{\top}$ gives
$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t)}-\alpha\bar{g}^{(t)}+\epsilon^{(t)},$
@@
where $\epsilon^{(t)}=\alpha n^{-1}(\mathds{1}^{\top}-\pi_{A}^{\top}D_{t}^{-1}){\mathbf{g}}^{(t)}$ decays exponentially fast [39, Lemma 12]. Setting $\hat{x}^{(t)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t)}$ recovers the form (7).
3.2 Push–Pull does not explicitly align with centralized SGD
Unlike Push-Only and Pull-Only, the Push-Pull method cannot be perfectly aligned with centralized SGD. Although it can be written in the form of (7), it suffers from a non-vanishing error term. Recall the Push-Pull recursion (2). Left-multiplying (2b) by $\mathds{1}^{\top}$ yields
$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(t+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(t)}+\mathds{1}^{\top}{\mathbf{g}}^{(t+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(t)}=\mathds{1}^{\top}{\mathbf{g}}^{(t+1)},$ (10)
where the last equality follows from the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying step (2a) by $\pi_{A}^{\top}$, we obtain
$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(t+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(t)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{(t)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}$
$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(t)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{(t)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}.$
Letting $\hat{x}^{(t)}=\pi_{A}^{\top}{\mathbf{x}}^{(t)}$ and using (10), we obtain
@@ 3.2 Push–Pull does not explicitly align with centralized SGD
$\displaystyle\hat{x}^{(t+1)}=\hat{x}^{(t)}-c\alpha\bar{g}^{(t)}+\epsilon^{(t)},\quad\text{where}\quad\epsilon^{(t)}=-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)},$ (11)
with $c=n\pi_{A}^{\top}\pi_{B}$ (see Definition (6)) and $\bar{g}^{(t)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(t)}$. Unlike Push-DiGing (Example 1), where $\epsilon^{(t)}=0$, or FROST (Example 2), where $\epsilon^{(t)}$ decays exponentially, the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}$ in Push-Pull does not vanish but instead stabilizes at a non-negligible constant. As shown in Figure 1, this persistent error remains regardless of network size or topology, making the linear speedup analysis of Push-Pull substantially more challenging.
Figure 1: With a fixed stepsize $\alpha$, the error term $\epsilon^{(t)}$ in Push-Pull does not vanish but rapidly stabilizes at a non-negligible constant. Results are shown for logistic regression on a synthetic dataset.
4 A Novel Multiple-Step Descent Analysis Framework
Establishing linear speedup for Push-Pull is challenging due to the presence of a non-vanishing error in (11). We begin by revisiting the classical single-step descent framework, highlighting why it suffices for Push-Only and Pull-Only methods but fails for Push-Pull, thereby motivating the development of a new multi-step descent analysis.
4.1 Limitations of single-step descent analysis
Following conventional single-step descent analysis, we establish the descent inequality based on update (11):
$\displaystyle\|\nabla f(\hat{x}^{(t)})\|^{2}\hskip-2.84526pt\lesssim\hskip-2.84526pt\frac{f(\hat{x}^{(t)})\hskip-2.84526pt-\hskip-2.84526pt\mathbb{E}_{t}[f(\hat{x}^{(t+1)})]}{c\alpha}\hskip-2.27621pt+\hskip-2.27621pt\frac{c\alpha L}{n}\sigma^{2}\hskip-2.27621pt+\hskip-2.27621pt\alpha L\|\pi_{A}^{\top}(I\hskip-2.84526pt-\hskip-2.84526ptB_{\infty}){\mathbf{y}}^{(t)}\|^{2}\hskip-2.84526pt+\hskip-2.84526pt\Delta_{\text{consent}}^{(t)}$ (12)
where $\Delta_{\text{consent}}^{(t)}$ denotes consensus errors among local variables, typically of order $\mathcal{O}(\alpha^{2})$ and hence negligible when $\alpha$ is sufficiently small. We use $\lesssim$ to denote inequalities that hold up to absolute numerical constants. Taking full expectation and summing over $t=0,\dots,T-1$ yields
$\displaystyle\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]$ $\displaystyle\lesssim\frac{f(\hat{x}^{(0)})\hskip-1.42262pt-\hskip-1.42262ptf(\hat{x}^{(T)})}{c\alpha T}\hskip-1.42262pt+\hskip-1.42262pt\frac{c\alpha L}{n}\sigma^{2}\hskip-1.42262pt\hskip-1.42262pt$ (13)
@@ 4.1 Limitations of single-step descent analysis
$\displaystyle\quad+\frac{\alpha L}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\pi_{A}^{\top}(I\hskip-1.42262pt-\hskip-1.42262ptB_{\infty}){\mathbf{y}}^{(t)}\|^{2}]+\frac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\Delta^{(t)}_{\rm consent}$]. (14)
Since $\tfrac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\Delta^{(t)}_{\rm consent}$] is negligible for sufficiently small $\alpha$, the convergence rate is dominated by the two main terms in (13), provided the first term in (14) is either zero (Example 1) or decays exponentially (Example 2). With $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$, we achieve $\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]=\mathcal{O}(\sigma/\sqrt{nT})$, implying $\mathcal{O}(1/(n\epsilon^{2}))$ iteration complexity and hence establish the linear speedup property.
However, unlike Push-Only and Pull-Only, the term $\mathbb{E}[\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}\|^{2}]$ in Push-Pull persists (see Figure 1). The central challenge in establishing linear speedup is therefore to prove the following inequality:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\pi_{A}^{\top}(I-B_{\infty})\mathbf{y}^{(t)}\|^{2}]\lesssim\frac{c}{n}\sigma^{2}+\text{negligible terms}.$ (15)
If this inequality holds, the first term in (14) can be absorbed into the second term of (13). Choosing $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$ will yield the same linear speedup rate. Unfortunately, this cannot be proved with standard techniques. Recall that $B_{\infty}=\pi_{B}\mathds{1}_{n}^{\top}$ and define $\Delta_{y}^{(t)}:=(I-B_{\infty})\mathbf{y}^{(t)}$. Left-multiplying (2b) by $(I-B_{\infty})$ gives
$\displaystyle\Delta_{y}^{(t+1)}=(B-B_{\infty})\Delta_{y}^{(t)}+(I-B_{\infty})\big({\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)}\big).$ (16)
where ${\mathbf{g}}^{(t)}$ stacks the $n$ stochastic gradients $\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})$ (each with variance $\sigma^{2}$), we can therefore derive that $\mathbb{E}[\|\pi_{A}^{\top}\Delta_{y}^{(t)}\|^{2}]=\Theta(n\sigma^{2})$, which is far larger than the desired $c\sigma^{2}/n$. This motivates the need for a fundamentally different approach.
4.2 A novel multi-step descent analysis
In contrast to analyzing each iteration in isolation, we now introduce a novel $m$-step recursion that aggregates information across a block of $m$ iterations. This blockwise perspective reveals cancellation effects in the gradient-tracking recursion and enables the gradient noise to be controlled at the desired level. Specifically, for a positive integer $m$ and any index $k$ that is a multiple of $m$, the recursion (11) can be written as
$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},$ (17)
where $\Delta_{y}^{(t)}=(I-B_{\infty})\mathbf{y}^{(t)}$. Analogous to the derivation of the single-step descent inequality (12), we can establish the following $m$-step descent inequality:
$\displaystyle\|\nabla f(\hat{x}^{(k)})\|^{2}\hskip-2.27621pt\lesssim\hskip-2.27621pt\frac{f(\hat{x}^{(k)})\hskip-2.27621pt-\hskip-2.27621pt\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]}{c\alpha m}\hskip-2.27621pt+\hskip-2.27621pt\frac{c\alpha L}{n}\sigma^{2}\hskip-2.27621pt+\hskip-2.27621pt\frac{\alpha L}{m}\|\pi_{A}^{\top}\hskip-2.27621pt\sum_{i=0}^{m-1}\hskip-2.84526pt\Delta_{y}^{(k+i)}\hskip-0.56905pt\|^{2}\hskip-2.27621pt+\hskip-2.27621pt\Delta_{\text{consent}}^{(k,m)}$ (18)
where $\Delta_{\text{consent}}^{(k,m)}$ denotes consensus errors within the $m$-step block, which are typically $\mathcal{O}(\alpha^{2})$ and negligible for small $\alpha$. A formal statement of (18) is provided in Lemma 3. Let $T=mK$ denote the total number of iterations and define $[m^{*}]:=\{0,m,2m,\ldots,(K-1)m\}$. Taking the full expectation of (18) and summing over all $k\in[m^{*}]$ yields
$\displaystyle\quad\min_{0\leq t\leq T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]$ $\displaystyle\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\lesssim\frac{f(\hat{x}^{(0)})-f(\hat{x}^{(T)})}{c\alpha T}+\frac{c\alpha L}{n}\sigma^{2}$
$\displaystyle\hskip-22.76228pt+\frac{\alpha L}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]+\frac{\alpha L}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{\text{consent}}^{(k,m)}].$ (19)
Similar to the single-step descent inequality (14), the central component in the $m$-step analysis is to control $\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}$. Summing the single-step recursion (16) from $i=0$ to $m-1$ yields
$\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\Delta_{y}^{(k)}+(B-B_{\infty})\sum_{i=0}^{m-2}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m-1)}-{\mathbf{g}}^{(k)}).$
This telescoping structure cancels all intermediate stochastic gradients ${\mathbf{g}}^{(k+i)}$ for $1\leq i\leq m-2$, leaving only boundary terms $\mathbf{g}^{(k)}$ and $\mathbf{g}^{(k+m-1)}$. Repeatedly applying this argument yields the equality
$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\Big(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\Big)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (20)
Unlike the single-step recursion (15), the $m$-step formulation cancels all interior noise, leaving only boundary differences ${\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}$, each attenuated by exponentially decaying factors $B^{m-j-1}-B_{\infty}$ (see Proposition 2). As a result, the block-level noise remains bounded at a constant level, independent of $m$.
In particular, each boundary difference contributes gradient noise with variance $n\sigma^{2}$, but its effect is scaled by $\lambda_{B}^{m-j-1}$, yielding $\sum_{j=1}^{m-1}\lambda_{B}^{m-j-1}n\sigma^{2}\leq(n\sigma^{2})/(1-\lambda_{B})$ which is constant in $m$. Therefore,
$\displaystyle\ \frac{1}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\pi_{A}^{\top}\Big(\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\Big)\|^{2}]$ (21)
$\displaystyle\lesssim$ $\displaystyle\ \frac{n\sigma^{2}}{m}+\text{negligible terms}\lesssim\frac{\sigma^{2}}{n}+\text{negligible terms}.$
The second inequality holds when $m$ is chosen sufficiently large [^1] In rigorous proofs, the choice of $m$ is constrained by additional factors and is therefore more involved than the simplified condition stated here, see more details in Section 5., e.g., $m\geq\text{Const}\cdot n^{2}$. Recalling that $T=mK$, substituting (21) into (18) and choosing $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$ yield the desired linear speedup rate. This highlights the unique advantage of the multiple-step descent analysis: by aggregating errors across $m$ steps within a block, it reveals a telescoping cancellation in the gradient-tracking recursion, leaving only boundary terms that are exponentially decayed during iterations. Consequently, the effective gradient noise is controlled at the desired $\mathcal{O}(\sigma^{2}/n)$ level, enabling the first proof of linear speedup for Push-Pull on strongly connected digraphs.